OpenStudy (anonymous):

Can you check my work on this problem? ( i accidentally copied the wrong problem. The new one is at the end)

OpenStudy (anonymous):

Is it 8 or is it no solution?

OpenStudy (unklerhaukus):

\[\sqrt{4x−7}−2=3\]you can always check solutions like this if x =8 is a solution \[\sqrt{4\times8−7}−2=3\]\[\sqrt{32−7}−2=3\]\[\sqrt{25}−2=3\]\[{5}−2=3\]\[3=3\]which is a true statement

OpenStudy (unklerhaukus):

so x=8 is a solution

OpenStudy (anonymous):

But what is the solution set then? 8,3?

OpenStudy (unklerhaukus):

what would the equation become if x=3?

OpenStudy (anonymous):

sqrt5-3 which is not 3. So how do i find the whole set? Or is it just eight?

OpenStudy (anonymous):

Oops. Sorry guys but i messed up on the problem. Its actually \[\sqrt{4x-7}-1=2\] Not really sure how i messed that one up. Thats what i get for doing math this late at night.

OpenStudy (unklerhaukus):

\[\sqrt{4x-7}−1=2\] add one to both sides \[\sqrt{4x-7}=3\]square both sides \[4x-7=9\] \[4x=16\] ...